# GUing: A Mobile GUI Search Engine using a Vision-Language Model

### v1

@@ Abstract.
App developers use the Graphical User Interface (GUI) of other apps as an important source of inspiration to design and improve their own apps. In recent years, research suggested various approaches to retrieve GUI designs that fit a certain text query from screenshot datasets acquired through automated GUI exploration. However, such text-to-GUI retrieval approaches only leverage the textual information of the GUI elements in the screenshots, neglecting visual information such as icons or background images. In addition, the retrieved screenshots are not steered by app developers and often lack important app features, e.g. whose UI pages require user authentication.
To overcome these limitations, this paper proposes GUing, a GUI search engine based on a vision-language model called UIClip, which we trained specifically for the app GUI domain. For this, we first collected app introduction images from Google Play, which usually display the most representative screenshots selected and often captioned (i.e. labeled) by app vendors. Then, we developed an automated pipeline to classify, crop, and extract the captions from these images. This finally results in a large dataset which we share with this paper: including 303k app screenshots, out of which 135k have captions. We used this dataset to train a novel vision-language model, which is, to the best of our knowledge, the first of its kind in GUI retrieval. We evaluated our approach on various datasets from related work and in manual experiment. The results demonstrate that our model outperforms previous approaches in text-to-GUI retrieval achieving a Recall@10 of up to 0.69 and a HIT@10 of 0.91. We also explored the performance of UIClip for other GUI tasks including GUI classification and Sketch-to-GUI retrieval with encouraging results.
@@ 1. Introduction
The Graphical User Interfaces (GUI) of mobile apps serve as the primary means of interaction between users and their devices. A well-designed GUI can markedly streamline the navigation process, facilitate the accomplishment of user tasks, and improve the overall user experience— contributing towards a higher user engagement and retention (Hassan et al., 2020; Chen et al., 2021). Furthermore, a user-friendly, modern, and attractive GUI design can potentially differentiate an app from its market counterparts, thereby amplifying its likelihood of success in the highly competitive mobile app market (Moran et al., 2018).
To design a good GUI, app designers and developers may relay on multiple sources. In addition to the initial needs and requirements of the users, designers often initiate the ideation process with inspiration searches of other GUIs. Numerous GUI retrieval approaches have thus been suggested, enabling the retrieval of GUIs from existing apps. Researchers have proposed GUI retrieval approaches that accept sketches (Huang et al., 2019; Mohian and Csallner, 2022, 2023), wireframes (Deka et al., 2017; Liu et al., 2018; Chen et al., 2020), or screenshots (Li et al., 2021; Bunian et al., 2021) as input to locate similar or fitting designs. While certainly useful, these approaches require a preliminary graphical prototype, which might not be available or might be to restrictive for the ideation in early development phases. Therefore, text-to-GUI retrieval approaches are more helpful when only ideas and requirements in text form are available.
Figure 1. Overview of our approach: including the creation of the datasets, vision-language model training, development of the GUI search engine, and the process of querying on the search engine.
@@
Recent GUI search engines, such as GUIGLE (Bernal-Cardenas et al., 2019) and RaWi (Kolthoff et al., 2023), enable users to search GUI datasets using text queries. Both models use of the app metadata and GUI-related text for query matching. GUIGLE adopts basic keyword matching, while RaWi calculates semantic similarity through a BERT model. However, these purely text-based approaches solely use textual information available with the GUI, such as text blocks or button labels, neglecting key visual information such as images, layout, or background. Furthermore, the GUI datasets used by these search engines (namely Rico (Deka et al., 2017) and ReDraw (Moran et al., 2020)) are created by automatically exploring app screenshots at runtime. However, the access to certain pages may require app authorization or initial configurations, which are often lacking in automatic exploration. Thus, screenshots scrawled at runtime may fail to comprehensively encapsulate the essential features of the app.
We propose a novel GUI search engine based on vision-language machine learning models and the app introduction images from Google Play as shown on Figure 1. Recent Vision-Language Models (VLMs), such as CLIP (Radford et al., 2021), BLIP (Li et al., 2022a), and BLIP-2 (Li et al., 2023), are trained in large-scale image-caption data by contrastive learning. These models have the ability to transform images and text into a multimodal embedding, ensuring that semantically similar images and texts are mapped closely in the embedding space. By computing the similarity between the images and the textual query, these models can be used in text-to-image retrieval tasks. For the effective application of the VLM in the app GUI domain, a large screenshot-caption dataset is a prerequisite. Currently available screenshot-caption datasets (Screen2Words (Wang et al., 2021) and Clarity (Moran et al., 2022)) are inadequate for training a VLM as our evaluation shows in Section 5. Therefore, we have created a new large dataset.
@@
Figure 2. Examples of app introduction images on Google Play.
Google Play is one of the most extensive mobile app stores. It offers thousands of apps from diverse domains, which makes it an abundant source of inspiration for requirements elicitation and app design (Ferrari and Spoletini, 2023). The app introduction images on Google Play are a gold mine for GUI retrieval, as they are carefully selected by app developers to represent the important features of the apps. Figure 2 shows examples of these images. A large portion of the app introduction images can be categorized as surrounded screenshots, each of them contains a screenshot as well as a caption that succinctly describe the screenshot, as shown on Figure 2 (a). In order to extract the screenshots and respective captions, we developed an automated pipeline, optimized for these specific images. The pipeline includes image classification (to classify image into three categories as shown in Figure 2), screen cropping (for cropping the screenshots area from the surrounded screenshots), and caption recognition (to extract captions from the surrounded screenshots, like the “LOG BLOOD PRESSURE” in Figure 2 (a)). We demonstrate the efficacy of our pipeline in Section 2. By applying the pipeline on app introduction images from approximately 117k apps, we created a comprehensive dataset comprising 303k screenshots, of which 135k have a caption. We refer to the collection of 303k screenshots as the GPSRepo (Google Play Screenshot Repository), and the subset containing captions as GPSCap (Google Play Screenshot Caption). By leveraging the GPSCap, in conjunction with Screen2Words and Clarity datasets, we fine tuned the CLIP model to create a VLM specific to the GUI domain. We call the new model UIClip. Evaluation results demonstrated a promising performance of our UIClip model for text-to-image retrieval. It superseded both the text-only approach and the CLIP model. Drawing upon the UIClip model and the GUIRepo dataset, we devised our GUI search engine, GUing. The engine accepts textual queries as input to retrieve relevant GUI images from our GPSRepo dataset as well as the Rico and ReDraw datasets. Our engine can easily be extended with additional screen datasets. Our empirical evaluations show that the search engine deliver highly relevant GUIs. Additional experiments suggest that the UIClip model is also beneficial for other GUI-related tasks, such as sketch-to-GUI retrieval and GUI classification. The paper provides the following contributions, with source code and datasets publicly available https://github.com/Jl-wei/guing:
  • A vision-language model named UIClip for a range of GUI-related tasks, including text-to-GUI retrieval, sketch-to-GUI retrieval, and GUI classification.
  • A GUI search engine that can achieve high performance with textual query.
@@ 2. Dataset Creation
In this section, we provide a comprehensive description of our datasets:
  • Google Play Screenshot Caption (GPSCap): This is a screenshot-caption dataset containing 135k pairs of screenshots and captions which serve as the training data for our vision language model.
  • Google Play Screenshot Repository (GPSRepo): This screenshot repository is composed of 303k screenshots which form the source of our search engine.
As depicted in Figure 3, the initial step involved collecting an extensive number of app introduction images from Google Play. Subsequently, we developed an image classifier to categorize the images into screenshots, surrounded screenshots and irrelevant. The area of surrounded screenshots containing screenshots were precisely cropped from the surrounding imagery through the application of object detection. Additionally, the captions from the surrounded screenshots were extracted with the aid of OCR (Optical Character Recognition). The images classified as screenshots and the screenshots cropped from surrounded screenshots constitute the GPSRepo, while the caption and screenshot pairs extracted from surrounded screenshots constitute the GPSCap. In the subsequent sections, we will describe the construction details of our datasets.
Figure 3. Overview of dataset creation pipeline
@@ 2.1. Image Collection
Given the ID of an app, we can easily download the app introduction images via Google Play Scraper (Olano, 2015). In order to obtain the app IDs, we initially gathered top ranked apps in each category from AppBrain (AppBrain, 2024a), creating a seed list. An app on Google Play often suggests similar apps, and a developer associated with the app may also offer additional apps. This scenario can be conceptually treated as a graph, where apps represent nodes and the relationships (such as app similarity and common developer) represent edges. Starting from the seed list, after performing breadth-first searching this graph, we collected a total of 117,283 apps (gathered on October 30, 2023). All apps were collected from Google Play in the United States in English language.
The stylistic differences between the introduction images of games and non-game apps render the former unsuitable as references for general app design. Consequently, we excluded games from our analysis, leaving a dataset of 85,631 non-game apps. These include a total of 874,130 images. However, a portion of these images were specifically demonstrated for landscape, Wear OS or Android TV (Google, 2024). To exclude them from our dataset, we applied a filtering process based on aspect ratio, leading to 553,755 precisely filtered images. Additionally, a subset of images could be duplicates despite bearing different filenames. In order to eliminate duplications, we implemented a process where we measured the SHA-1 hash of each individual image, retaining only one image with the same hash. This process yielded a final count of 398,511 unique app introduction images.
@@ 2.2. Image Classification
The images obtained from the prior phase are classified into three categories for subsequent analysis, as depicted in Figure 3. A subset of these images encompasses the complete display of an app user interface: those are screenshots. For some images, only a segment of the image represents a captured screenshot, and are thus categorized as surrounded screenshots. The remaining images do not encapsulate screenshot or only contain partial or slanted screenshot are accordingly classified as irrelevant. Manual classification on around 400k images is laborious and time-consuming. Thus, we trained an image classifier to classify automatically these images into three categories: screenshots, surrounded screenshot and irrelevant.
@@ 2.2.1. Classification Model
The Vision Transformer (ViT) (Dosovitskiy et al., 2021) was employed as the classifier for our dataset, based on its performance metrics, which will be discussed in subsequent sections. The ViT architecture essentially constitutes the encoder of transformer as described by Vaswani et al. (Vaswani et al., 2017). It has demonstrated superior performance in image classification tasks. Within the realm of natural language processing, the dominant approach involves pre-training the model on a large text corpus followed by fine-tuning on a more targeted dataset (Devlin et al., 2019). Similarly, the ViT benefits from a pre-training phase using a comprehensive collection of images, thereby allowing the model to capture innate image features. Consequently, when deployed on novel tasks, the ViT requires a reduced amount of task-specific training due to the knowledge acquired during this pre-training process.
2.2.2. Training Data Creation
To train and fine tune the classification model, a labeled dataset is required. For this, we randomly sampled 3,615 images from the filtered images collected on Section 2.1. To annotate these images, we created an annotation guide to clarify the definition of screenshots, surrounded screenshot and irrelevant. Two of the authors then annotated these images independently, the conflicting labels were reviewed through discussion and consensus.
@@ 2.2.3. Classification Performance
The performance of the classifier was evaluated using the app introduction images from the dataset annotated in the previous step, which was split in an 80:20 ratio for training and testing respectively. The classification performance was measured by precision, recall and F1. We trained the classification model using mini-batch gradient descent, with a batch size of 64 and AdamW optimizer with an initial learning rate of $2e^{-5}$. It was trained on 5 epochs on a machine with a NVIDIA Tesla T4 GPU with 16 GB VRAM. In order to enhance the robustness of our results, we performed ten-fold cross-validation by repetitively splitting the dataset into distinct training and testing subsets ten times in a randomized manner. We then computed the mean performance across these runs. The experiment results presented in Table 1 show the great performance of our classifier, with an average F1 score 0.964.
Table 1. Image classification accuracy on evaluation set
Precision Recall F1
@@
2.2.4. Applying Classifier
The evaluation demonstrated the robust performance of our classifier. Subsequent to that evaluation, we trained a ViT model on all of the 3,615 images. This model was then employed to classify all of the filtered app introduction images. To augment the reliability and integrity of our GUI repository, we calibrated the classifier’s threshold at a high confidence level of 0.9, which means that only images that attain a classification probability exceeding the 0.9 are included in the respective categories. Our empirical investigations reveal that, at the 0.9 threshold, the precision scores for the categories screenshots and surrounded screenshots surpass 0.99. Following the classification process, our repository had 111,847 images classified as screenshots and 191,993 as surrounded screenshots.
@@ 2.3. Screen Cropping
The surrounded screenshots are the images that contain screenshots with additional visual frames. To automatically crop the screenshot areas from these surrounded screenshots, we trained an object detection model. This model is capable of localizing the screenshot within the larger image and subsequently generating a bounding box. A bounding box is defined as a rectangular delineation that encases the area of interest: in this case, the screenshot. Given the bounding boxes inside the surrounded screenshots, one can easily crop the screenshot area using an image processing library.

### v2

@@ Abstract.
App developers use the Graphical User Interface (GUI) of other apps as a source of inspiration for designing and improving their own apps. Recent research has thus suggested retrieving relevant GUI designs that match a certain text query from screenshot datasets acquired through crowdsourced or automated exploration of GUIs. However, such text-to-GUI retrieval approaches only leverage the textual information of the GUI elements, neglecting visual information such as icons or background images. In addition, retrieved screenshots are not steered by app developers and often lack important app features that require particular input data.
To overcome these limitations, this paper proposes GUing, a GUI search engine based on a vision-language model called GUIClip, which we trained specifically for the problem of designing app GUIs. For this, we first collected from Google Play app introduction images which usually display the most representative screenshots and are often captioned (i.e. labeled) by app vendors. Then, we developed an automated pipeline to classify, crop, and extract the captions from these images. This resulted in a large dataset which we share with this paper: including 303k app screenshots, out of which 135k have captions. We used this dataset to train a novel vision-language model, which is, to the best of our knowledge, the first of its kind in GUI retrieval. We evaluated our approach on various datasets from related work and in manual experiment. The results demonstrate that our model outperforms previous approaches in text-to-GUI retrieval achieving a Recall@10 of up to 0.69 and a HIT@10 of 0.91. We also explored the performance of GUIClip for other GUI tasks including GUI classification and sketch-to-GUI retrieval with encouraging results.
ccs: Human-centered computing Graphical user interfacesccs: Software and its engineering Designing software
@@ 1. Introduction
The Graphical User Interfaces (GUI) of mobile apps serve as the primary means of interaction between users and their devices. A well-designed GUI streamlines the navigation process, facilitates the accomplishment of user tasks, and improves the overall user experience— contributing towards a higher user engagement and retention (Hassan et al., 2020; Chen et al., 2021). Furthermore, a modern, attractive, and user-friendly GUI can potentially differentiate an app from its market counterparts, amplifying its likelihood of success in the highly competitive app market (Moran et al., 2018; Martens and Maalej, 2019b).
To design a good GUI, app developers may rely on multiple sources. In addition to initial requirements elicitation and prototyping, app developers often explore other GUIs of related apps. Numerous approaches have thus been suggested to retrieve relevant GUIs from existing apps. Researchers have proposed GUI retrieval approaches that accept sketches (Huang et al., 2019; Mohian and Csallner, 2022, 2023), wireframes (Deka et al., 2017; Liu et al., 2018; Chen et al., 2020), or screenshots (Li et al., 2021; Bunian et al., 2021) as input to locate similar or fitting designs. While certainly useful, these approaches require a preliminary graphical prototype, which might not be available or might be too restrictive in early ideation phases. Therefore, when only general ideas and textual requirements are available, text-to-GUI retrieval approaches would be particularly useful.
Figure 1. Overview of our approach: including the creation of the datasets, the training of the vision-language model GUIClip, the development of the GUI search engine, and the process of querying on the engine.
@@
Recent GUI search engines, such as GUIGLE (Bernal-Cardenas et al., 2019) and RaWi (Kolthoff et al., 2023), enable their users to search GUI datasets using text queries. Both engines use app metadata and GUI-related text for query matching. GUIGLE adopts basic keyword matching, while RaWi calculates semantic similarity based on a BERT model (Devlin et al., 2019). However, these text-based retrieval approaches solely use textual information available within the GUI such as text blocks or button labels, neglecting key visual information such as images, layouts, and backgrounds. Furthermore, the underlying GUI datasets (namely Rico (Deka et al., 2017) and ReDraw (Moran et al., 2020)) are created by crowdsourced or automated exploration of app screens at runtime. However, the access to certain screens may require authorization or initial configurations, which is often skipped in automated exploration. Even with a crowdsourced exploration, some app features may not be captured, particularly those requiring substantial or specific input data, such as dashboard pages with charts. Thus, screenshots crawled at runtime may fail to comprehensively capture essential features of the app.
In this paper, we propose a novel GUI search engine based on vision-language machine learning models trained with app introduction images from Google Play as shown on Figure 1. Recent Vision-Language Models (VLMs), such as CLIP (Radford et al., 2021), BLIP (Li et al., 2022a), and BLIP-2 (Li et al., 2023), are trained on large-scale image-caption data using contrastive learning. These models have the ability to transform images and text into a multimodal embedding, ensuring that semantically similar images and texts are mapped closely in the embedding space. Thus, by computing the similarity between the images and the textual query, these models can be used for text-to-image retrieval tasks. As prerequisite a large screenshot-caption dataset is needed. However, currently available datasets (Screen2Words (Wang et al., 2021) and Clarity (Moran et al., 2022)) are inadequate for training a VLM as our evaluation shows in Section 5. Therefore, we have created a new large screenshot-caption dataset.
@@
Figure 2. Examples of app introduction images from Google Play app store.
Google Play is one of the largest mobile app stores with thousands of apps from diverse domains. This makes it an rich source of inspiration for requirements elicitation and app design (Ferrari and Spoletini, 2023; Maalej et al., 2024). Particularly, the app introduction images on Google Play are a gold mine for GUI retrieval, as they are carefully selected and described by app developers to represent the important app features. Figure 2 shows examples of these images. A large portion can be categorized as surrounded screenshots, each displaying a screenshot and a succinct caption that describes the screen (Figure 2 (a)). In order to extract the screenshots and respective captions, we developed an automated pipeline, optimized for these images. The pipeline first classifies images into three categories as shown in Figure 2. Then it crops the screenshots area from the surrounded screenshots and extracts the caption from the surrounded screenshots (like the “LOG BLOOD PRESSURE” in Figure 2 (a)). We demonstrate the efficacy of our pipeline in Section 2.
By applying the pipeline on the introduction images of approximately 117k apps, we created a comprehensive dataset comprising 303k screenshots, of which 135k have captions. We refer to the collection of 303k screenshots as the ScreenRepo and the subset containing captions as SCapRepo. We then used SCapRepo together with the Screen2Words and Clarity datasets to fine-tune the CLIP model and create a VLM specific to the GUI domain. We call the new model GUIClip. Our evaluation results demonstrate a promising performance of GUIClip for text-to-image GUI retrieval, outperforming both text-only approaches and the CLIP model. Based on GUIClip, we developed GUing, a search engine, that accepts textual queries as input to retrieve relevant GUI images from our ScreenRepo dataset as well as the Rico and ReDraw datasets. Our engine can easily be extended with additional screen datasets. Our manual evaluation shows that the search engine recommends highly relevant GUIs. Additional experiments suggest that GUIClip is also beneficial for other GUI-related tasks, such as sketch-to-GUI retrieval and GUI classification. The paper provides the following contributions, with source code and datasets publicly available https://github.com/Jl-wei/guing:
  • A vision-language model named GUIClip for a range of GUI-related tasks, including text-to-GUI retrieval, sketch-to-GUI retrieval, and GUI classification available at https://huggingface.co/Jl-wei/guiclip-vit-base-patch32.
  • A GUI search engine that can achieve high performance with textual queries.
@@ 2. Dataset Creation
Created two datasets: (1) Google Play Screenshot Caption (SCapRepo) including 135k screenshot-caption pairs for the training of our vision-language model and (2) Google Play Screenshot Repository (ScreenRepo) including 303k screenshots for the search engine repository. Figure 3 depicts the datasets creation pipeline. In the initial step, we collected app introduction images from Google Play. Subsequently, we developed an image classifier to categorize the images into screenshots, surrounded screenshots and irrelevant. The area in a surrounded screenshot containing a screenshots was precisely cropped from the surrounding by applying object detection. Additionally, we extracted the captions from the surrounded screenshots by applying Optical Character Recognition (OCR). The images classified as screenshots together with the cropped screenshots constitute the ScreenRepo, while the screenshot-caption pairs extracted from surrounded screenshots constitute the SCapRepo. The following describes each step in detail.
Figure 3. Overview of dataset creation pipeline.
@@ 2.1. Image Collection
Given the ID of an app, we can download the app introduction images via Google Play Scraper (Olano, 2015). In order to obtain the app IDs, we initially gathered top ranked apps in each category from AppBrain (AppBrain, 2024), creating a seed list. An app on Google Play often suggests similar apps and the app developer who may also offer additional apps. This scenario can be conceptually treated as a graph, where apps represent nodes and the relationships (such as app similarity and common developer) represent edges. Starting from the seed list, we performed a breadth-first search on this graph, collecting a total of 117,283 apps (by October 2023). All apps were collected from the US Google Play in English language.
The stylistic differences between the introduction images of games and non-game apps make the former unsuitable as reference for general app design. Consequently, we excluded games from our analysis, leaving a dataset of 85,631 non-game apps. This includes a total of 874,130 images. Some of these images specifically demonstrates landscape views, Wear OS, or Android TV (Google, 2024). To exclude those from our dataset, we applied a filtering process based on aspect ratio, leading to 553,755 precisely filtered images. Additionally, some images could be duplicates despite having different filenames. To remove duplicates, we calculated the SHA-1 hash of each individual image, retaining only one image with the same hash. This yielded a final count of 398,511 unique app introduction images.
@@ 2.2. Image Classification
For subsequent analysis, we classified the images obtained from the prior phase into three categories as depicted in Figure 3. A subset of these images show the complete display of an app user interface: those are screenshots. In some images, only a segment represents a captured screenshot. Those are categorized as surrounded screenshots. The remaining images either do not show a screenshot or only a partial or a slanted screenshot. Those are classified as irrelevant. As a manual classification of around 400k images is unreasonable, we trained an image classifier to automatically classify these images into screenshots, surrounded screenshot, or irrelevant.
@@ 2.2.1. Classification Model
We used the Vision Transformer (ViT) (Dosovitskiy et al., 2021) as classifier due to its top performance (as will be discussed in subsequent sections). The ViT architecture essentially constitutes the encoder of transformer as described by Vaswani et al. (Vaswani et al., 2017). It has demonstrated superior performance in image classification tasks. In the area of natural language processing, the dominant approach is to pre-train the model on a large text corpus followed by fine-tuning on a more targeted dataset (Han et al., 2021). Similarly, the ViT benefits from a pre-training phase using a comprehensive collection of images, allowing the model to capture inherent image features. Consequently, when deployed on novel tasks, the ViT requires a reduced amount of task-specific training due to the knowledge acquired during its pre-training.
2.2.2. Creation of Training Data
To train and fine-tune the classification model, a labelled dataset is required. For this, we randomly sampled 3,615 images from the collected filtered dataset (see Section 2.1). To annotate these images, we created an annotation guide with examples to clarify the definition of screenshots, surrounded screenshot, and irrelevant. Two of the authors then annotated these images independently. For the entire labelled dataset, only 7 images had conflicting labels, primarily due to disagreements between the surrounded screenshot and irrelevant categories. These conflicts arose because the images in question did not contain a full screenshot but rather a UI component, such as a widget. After discussion, we reached a consensus to label images featuring a larger UI component (as thus potentially inspiring for a screen design) as surrounded screenshot, and the others as irrelevant.
@@ 2.2.3. Classification Performance
We evaluated the performance of the classifier using the labelled app introduction images from the previous step, which we split into a 80:20 ratio for training and testing, respectively. We trained the classification model using mini-batch gradient descent, with a batch size of 64 and AdamW optimiser with an initial learning rate of $2e^{-5}$. The classifier was trained in 5 epochs on a machine with a NVIDIA Tesla T4 GPU with 16 GB VRAM. In order to enhance the robustness of our results, we performed a 10-fold cross-validation by repetitively splitting the dataset into distinct training and testing subsets ten times in a randomized manner. We then computed the mean values of the precision, recall, and F1 score across these runs. The results are presented in Table 1 and show a top performance of our classifier, with an average F1 score of 0.964.
Table 1. Accuracy of the image classification (mean values for 10-fold cross-validation).
Class Precision Recall F1
@@
2.2.4. Applying the Classifier
Subsequent to the evaluation, we trained a ViT model on all of the 3,615 images. This model was then used to classify all of the filtered app introduction images in the dataset. To increase the reliability and integrity of our repository, we calibrated the classifier’s threshold at a high confidence level of 0.9, which means that only images with a classification probability exceeding 0.9 are included in the respective category. Our empirical investigation reveals that, at the 0.9 threshold, the precision scores for the categories screenshots and surrounded screenshots surpass 0.99. Finally, our repository had 111,847 images classified as screenshots and 191,993 as surrounded screenshots.
@@ 2.3. Screen Cropping
The surrounded screenshots are the images that contain screenshots with additional visual frames. To automatically crop the screenshot areas from these surrounded screenshots, we trained an object detection model. This model is capable of localizing the screenshot within the larger image and subsequently generating a bounding box. A bounding box is defined as a rectangular delineation that encases the area of interest: in this case, the screenshot. Given the bounding boxes inside the surrounded screenshots, we can easily crop the screenshot area using an image processing library.